# Proportionality for Constrained Public Decisions

### v1

@@ Abstract
We study situations where a group of voters need to take a collective decision over a number of public issues, with the goal of getting a result that reflects the voters’ opinions in a proportional manner. Our focus is on interconnected public decisions, where the decision on one or more issues has repercussions on the acceptance or rejection of other public issues in the agenda. We show that the adaptations of classical justified-representation axioms to this enriched setting are always satisfiable only for restricted classes of public agendas. However, the use of suitably adapted well-known decision rules on a class of quite expressive constraints, yields proportionality guarantees that match these justified-representation properties in an approximate sense. We also identify another path to achieving proportionality via an adaptation of the notion of priceability.
@@ 1 Introduction
In many situations of collective decision-making, a group of voters is presented with a set of issues for which they are expected to make a binary choice: typically, deciding to either accept or reject each issue. This setting has recently been studied under the name of public decisions (Skowron and Górecki, 2022) and it is of particular interest due to the real-world scenarios captured by it. Notable examples include: instances of multiple referenda where the public vote directly on the resolution of political issues; group activity planning, where a group of individuals are to choose, as a collective, the activities that the entire group shall partake in; and committee elections, where a set of candidates are in the running for multiple positions on a committee and a group of decision-makers must select the committee members (Lang and Xia, 2016).
@@
Given the collective nature of the problem, one of the natural desiderata is that the outcome represents a fair compromise for the participating voters. Among the numerous possible interpretations of fairness is the one captured by the notion of proportional representation. Proportionality features prominently in many collective choice settings such as that of apportionment (Balinski, 2005) and the aforementioned committee elections (Lackner and Skowron, 2023) while being introduced to richer social-choice models such as that of participatory budgeting (PB) (Rey and Maly, 2023). Indeed, even when zooming in on the public decisions task, the goal of producing collective outcomes that proportionally reflect the opinions of the voter population has been drawing increasing attention in recent years (Freeman et al., 2020; Masařík et al., 2023; Skowron and Górecki, 2022). However, a component that has so far not received much attention in this growing literature on proportionality is the presence of constraints that restrict the possible outcomes that can be returned. In this paper, we focus on answering the question of what one may do when outcomes that would satisfy classical proportionality axioms—and thus be considered fair outcomes—are no longer feasible due to the presence of constraints.
When examining real-world examples of the public-decision model, there are many scenarios where enriching the model with constraints fits naturally: in the case of participatory budgeting, the implementation of one project may be conditional on the acceptance (or rejection) of another; diversity constraints applied to the committee election problem that determine the number of individuals with certain characteristics that may be accepted/rejected; or when selecting the features of some product, only certain feature combinations represent affordable options.
In tackling our task, we build on existing notions of proportionality that have been posed for less rich models and tailor them for the challenges that comes with the existence of constraints. Naturally, this leads us to also consider constrained versions of collective decision rules proposed in the literature and to investigate the extent to which they meet the requirements of our novel constraint-aware notions of proportionality.
Related work. We begin by noting that our constrained public-decision model closely resembles that of judgment aggregation and it also naturally fits into the area of collective decisions in combinatorial domains (see (Endriss, 2016) and (Lang and Xia, 2016) for general introductions to these two topics, respectively).
Most relevant to our paper is the recent work conducted on fairness in the context of public decisions without constraints (Conitzer et al., 2017; Freeman et al., 2020; Masařík et al., 2023; Skowron and Górecki, 2022). Conitzer et al. (2017) focused on individually proportional outcomes, thus, our work more closely aligns with that of Freeman et al. (2020) and Skowron and Górecki (2022) who adapt the notion of justified representation (Aziz et al., 2017; Fernández et al., 2017; Peters and Skowron, 2020) from the literature of multiwinner voting (MWV) (Lackner and Skowron, 2023). Moreover, proportionality has also been studied in models of sequential decision-making that are relevant to our own as they can be seen as generalisations of the public-decision model without constraints (Bulteau et al., 2021; Chandak et al., 2024; Lackner, 2020). Amongst these sequential decision-making papers, those of Bulteau et al. (2021) and Chandak et al. (2024) relate to our work the most as they also implement justified-representation notions. More recently, Masařík et al. (2023) studied proportionality for a general social-choice model that allows for the modelling of both the unconstrained and constrained versions of the public-decision model. By focusing on the latter, we explore properties that are specifically made for this setting, which in turn allows us to define, and subsequently conduct an analysis of, constrained public-decision rules that are not touched upon by Masařík et al. (2023). Thus, our results complement their work by showing further possibilities, and also limitations, for proportionality within this constrained public-decision model. We also highlight work by Mavrov et al. (2023) who adapted justified representation for the MWV model with arbitrary constraints instead of our focus on the public-decision model. This leads us towards differing approaches in adapting justified representation for constraints and also, analysing quite different rules.
In related fields, previous work studied proportionality in various models that differ from the constrained public-decision model but features collective choices on interconnected propositions: the belief merging setting (Haret et al., 2020), interdependent binary issues via conditional ballots (Brill et al., 2023), and approval-based shortlisting with constraints (presented in a model of judgment aggregation) (Chingoma et al., 2022).
Contribution. We study the extent to which proportionality can be ensured constrained public-decision setting. First, we introduce the notion of feasible group deviations as a building block that allows the translation of existing proportionality axioms—that are based on varying public-decision interpretations of justified representation—for this setting with constraints.
@@
For each of our axioms, we show that although it is challenging to satisfy these properties in general constrained instances, when one hones in on a restricted—yet highly expressive—class of constraints, we can achieve proportionality guarantees that represent approximations of the desirable justified-representation axioms. In doing so, we also define novel adaptations of recently studied decision rules to our public-decision setting with constraints, namely the method of equal shares (MES) and the MeCorA rule. Finally, we adapt the priceability notion from the MWV literature, which provides another promising route to introduce proportionality into public decisions under constraints.
@@
Paper outline. We begin by detailing the constrained public-decision model in Section 2. We continue with Section 3 where we discuss two known ways in which justified representation is formalised for public decisions, and also present the notion of deviating groups. Then each of sections 4 and 5 deal with a particular public-decision interpretation of justified representation. Before concluding in Section 7, we deal wit our constrained version of the priceability axiom in Section 6. Note that all omitted proofs can be found as part of the supplementary material.
@@ 2 The Model
A finite set of $n$ voters $N=\{1,\ldots,n\}$ has to take a collective decision on a finite set of $m$ binary issues =$\mathcal{I}=\{a_{1},\ldots,a_{m}\}$. It is typical in the public decisions setting to consider there only being two available decisions per issue but we instead adopt the following, more general setup. Each issue $a_{t}\in\mathcal{I}$ is associated with a finite set of alternatives called a domain $D_{t}=\{d_{t}^{1},d_{t}^{2},\ldots\}\subseteq X$ where $|D_{t}|\geqslant 2$ holds for all $t\in[m]$. The design decision of going beyond binary issues is motivated by the wider real-life applicability of this model when more than two alternatives are possible for each issue.
Each voter $i\in N$ submits a ballot $\boldsymbol{b}_{i}=(\boldsymbol{b}_{i}^{1},\ldots,\boldsymbol{b}_{i}^{m})\in D_{1}\times\ldots\times D_{m}$ where $\boldsymbol{b}_{i}^{t}=d_{t}^{c}$ indicates that voter $i$ chooses the decision $d_{t}^{c}$ for the issue issue $a_{t}$.
A profile $\boldsymbol{B}=(\boldsymbol{b}_{1},\ldots,\boldsymbol{b}_{n})\in(D_{1}\times\ldots\times D_{m})^{n}$ is a vector of the $n$ voters’ ballots. An outcome $\boldsymbol{w}=(w_{1},\ldots,w_{m})\in D_{1}\times\ldots\times D_{m}$ is then a vector providing a decision for every issue at stake.
@@
We focus on situations where some constraints limit the set of possible collective outcomes: we denote by $\mathcal{C}\subseteq D_{1}\times\ldots\times D_{m}$ the set of feasible outcomes. We write $(\boldsymbol{B},\mathcal{C})$ to denote an election instance. By a slight abuse of notation we also refer to $\mathcal{C}$ as the constraint, and thus, we refer to elections instances where $\mathcal{C}=D_{1}\times\ldots\times D_{m}$ as unconstrained election instances.  [^1] Note that while we work formally with the constraint being an enumeration of all feasible outcomes, in practice, it is often possible to represent the set of feasible outcomes in more concise forms—via the use of formulas of propositional logic, for example—to help with parsing said constraint and/or speed up computation by exploiting the constraint’s representation structure.
Note that voter ballots need not be consistent with the constraints, i.e., for an election instance $(\boldsymbol{B},\mathcal{C})$, we do not require that $\boldsymbol{b}_{i}\in\mathcal{C}$ for all voters $i\in N$ [^2] This assumption takes our model closer to the particular model of judgment aggregation where the constraints on the output may differ from the constraints imposed on the the voters’ input judgments (Endriss, 2018; Chingoma et al., 2022).
Remark 1.
While not common in work done in the related judgment aggregation model, our assumption that voters ballots need not correspond to feasible outcomes is common in other settings of social choice. In multiwinner voting, voters can approve more candidates than the committee target size while in participatory budgeting, the sum of the costs of a voter’s approved projects may exceed the instance’s budget. For our setting, we argue that this approach helps capture real-world, constrained decision-making scenarios where either the constraint is uncertain when voters submit their ballots, or possibly, the voting process becomes more burdensome for voters as they attempt to create ballots with respect to a (possibly difficult to understand) constraint. For example, consider a group of friends deciding on the travel destinations of their shared holiday across the world, visiting one country in each continent. On a booking platform, there are a certain number of locations that can be selected per continent such as: Amsterdam, Paris and Vienna in Europe; Mexico City and Toronto in North America; Cairo, Nairobi and Cape Town in Africa; and so on. Each friend has a preferred combination of cities and their collective itinerary is subject to factors such as their travel budget or the available flight connections between cities. However, as flight costs and connections may change significantly on a day-to-day basis, it may be unclear which combination of cities are affordable. Therefore, it is not reasonable to impose the requirement, by default that is, that voter ballots are constraint-consistent.
If needed, we explicitly state when we pivot from this assumption and require that voter ballots be constraint-consistent. At times, we shall restrict ourselves to election instances where $D_{t}=\{0,1\}$ holds for every issue $a_{t}$. We refer to such cases as binary election instances. When necessary, we explicitly state whether any result hinges on the restriction to binary instances. Given an outcome $\boldsymbol{w}$ for a binary instance, the vector $\boldsymbol{\bar{w}}=({\bar{w}}_{1},\ldots,{\bar{w}}_{m})$ is such that $\bar{w}_{t}=1-w_{t}$ for all issues $a_{t}\in\mathcal{I}$.
Now, consider an outcome $\boldsymbol{w}$, a set of issues $S\subseteq\mathcal{I}$ and some vector $\boldsymbol{v}=(1,\ldots,v_{m})\in D_{1}\times\ldots\times D_{m}$ (that can be interpreted as either an outcome or voter’s ballot). We write ${\boldsymbol{w}}[{S\leftarrow\boldsymbol{v}}]=(w_{1}^{\prime},\ldots,w_{m}^{\prime})$ where $w_{t}^{\prime}=w_{t}$ for all issues $a_{t}\in\mathcal{I}\setminus{S}$ and $w_{t}^{\prime}=v_{t}$ for all issues $a_{j}\in S$. In other words, ${\boldsymbol{w}}[{S\leftarrow\boldsymbol{v}}]$ is the resultant vector of updating outcome $\boldsymbol{w}$’s decisions on the issues in $S$ by fixing them to those of vector $\boldsymbol{v}$. For a given issue $a_{t}\in\mathcal{I}$ and a decision $d\in D_{t}$, we use $N(a_{t},d)=\{i\in N\mid\boldsymbol{b}_{i}^{t}=d\}$ to denote the set of voters that agree with decision $d$ on issue $a_{t}$. Given two vectors $\boldsymbol{v},\boldsymbol{v}^{\prime}\in D_{1}\times\ldots\times D_{m}$, we denote the agreement between them by $\text{\rm Agr}(\boldsymbol{v},\boldsymbol{v}^{\prime})=\{a_{t}\in\mathcal{I}\mid v_{t}={v}_{t}^{\prime}\}$. Then, the satisfaction that a voter $i$ obtains from an outcome $\boldsymbol{w}$ corresponds to $u_{i}(\boldsymbol{w})=|\text{\rm Agr}(\boldsymbol{b}_{i},\boldsymbol{w})|$, i.e., the number of decisions on which the voter $i$ is in agreement with outcome $\boldsymbol{w}$.
3 Proportionality via Justified Representation
This section starts with the observation that classical notions of proportionality fall short when considering interconnected decisions (in the upcoming Example 1), and then follows with our proposed generalisations of such axioms that deal with constraints.
Ideally, when looking to make a proportional collective choice, we would like to meet the following criteria: a group of similarly-minded voters that is an $\alpha$ fraction of the population should have their opinions reflected in an $\alpha$ fraction of the $m$ issues. We wish to define an axiom for our model that captures this idea within our richer framework.
In the setting of multiwinner voting, this is formally captured with the justified representation axioms with one of the most widely studied being extended justified representation (EJR) (Aziz et al., 2017). Now, when being studied in the setting of public decisions, there are two different adaptations that have been studied and we shall look at both. One approach intuitively states that ‘a group of voters that agree on a set of issues $T$ and represent an $\alpha$ fraction of the voter population, should control a $\alpha\cdot|T|$ number of the total issues in $\mathcal{I}$(Chandak et al., 2024; Masařík et al., 2023; Skowron and Górecki, 2022). We refer to it as agreement-EJR.
This approach differs from the following that is a more faithful translation of the EJR from multiwinner voting: ‘a group of voters that agree on, and represent, an $\alpha$ fraction of the issues, and voter population, respectively, should control $\alpha\cdot m$ of the issues in $\mathcal{I}$(Chandak et al., 2024; Freeman et al., 2020). The requirements on the voter groups that is present in the latter approach are captured by the notion of cohesiveness and so we refer to this version of EJR as cohesiveness-EJR. Observe that cohesiveness-EJR is stronger than, and implies, agreement-EJR.
Meeting the ideal outlined by both of these notions is not easy in our setting as the constraint $\mathcal{C}$ could rule out a seemingly fair outcome from the onset.

### v2

@@ Abstract
We study situations where a group of voters need to take a collective decision over a number of public issues, with the goal of getting a result that reflects the voters’ opinions in a proportional manner. Our focus is on interconnected public decisions, where the outcome on one or more issues has repercussions on the acceptance or rejection of other issues in the agenda. We show that the adaptation of classical justified-representation axioms to this enriched setting are always satisfiable only for restricted classes of public agendas. We adapt well-known proportional decision rules to take the structure of the public agenda into account, and we show that they match justified-representation properties in approximation on a class of expressive constraints. We also identify another path to achieving proportionality on interconnected issues via an adaptation of the notion of priceability.
@@ 1 Introduction
Consider a municipality that is planning the renovation of one of its main squares in the historical centre. A project of this calibre implies taking numerous decisions, such as on whether to plant some trees (and of which kind), whether to add a fountain, some benches and tables, a bike-sharing station, a statue (and from which artist), and so on. Some of the potential configurations of the square may also be unfeasible, either because they would go over the allocated budget for the project, or because there may not be enough physical space to accommodate all of the desired features.
Given the impact that such a renovation would have on the everyday lives of its citizens, the municipality may decide to set up a participatory design process, in order to make sure that different preferences and needs are taken into account in the final design. [^1] Our example evokes the participatory process that was set up by the Municipality of Madrid in order to collectively redesign Plaza de España—although the project was later criticized by the opposition due to its low participation rate (Lyne, 2017). This is just one of many real-world examples of public decisions where (i) the final outcome should strive to be representative of the views of the different stakeholders involved, while also (ii) having some external constraints that rule out some possible outcomes as unfeasible.
Regarding our point (ii) above, indeed, the square re-design problem can be seen as a (generalized) participatory budgeting instance (cf. the recent survey by Rey and Maly, 2023), where the implementation of one project may be conditional on the acceptance, or rejection, of another, all the while satisfying some monetary and spatial constraints. We can also think of the collective product configuration problem (Astesana et al., 2010), where a group of people may need to collectively choose the features of a given product (e.g, a group of friends booking an accommodation for their holidays), knowing that some value combinations may be unavailable or out of stock. Finally, in a committee election (Lackner and Skowron, 2023) we may need to fill the positions in a committee with some candidates by also respecting various diversity constraints.
The latter example is an instance of an important sub-setting of public decisions, namely one where the issues at stake all have a binary domain: namely, the decision-makers are asked to express their preferences as an acceptance or a rejection of each issues (or candidate). These kind of scenarios may occur also in the context of multiple referenda, where the public votes directly on the resolution of political issues, and for group activity planning, where a group of individuals has to choose, as a collective, the activities that the entire group shall partake in (Lang and Xia, 2016).
@@
If we now turn to our point (i) above, among the numerous interpretations of fairness that have been put forward to make sure that the outcome appropriately reflect the diverse views of the participants, we will focus on the notion of proportional representation. Proportionality features prominently in many collective choice settings such as that of apportionment (Balinski, 2005) and the aforementioned committee elections (Lackner and Skowron, 2023). Indeed, even when zooming in on the problem of public decisions (Skowron and Górecki, 2022), the goal of producing collective outcomes that proportionally reflect the opinions of the voter population has been drawing increasing attention in recent years (Freeman et al., 2020; Masarík et al., 2024). However, a component that has so far not received much attention in this growing literature on proportionality is the presence of constraints that restrict the possible outcomes that can be returned, i.e., the combination of our points (i) and (ii) above. In this paper, we thus focus on answering the question of what one may do when outcomes that would satisfy classical proportionality axioms—and thus be considered fair—are no longer feasible due to the presence of constraints.
@@
Our contributions chart the extent to which proportionality can be ensured in a constrained public-decision setting. First, we use the notion of feasible group deviations as a building block that allows the adaptation of existing proportionality axioms—that are based on varying public-decision interpretations of justified representation—to decisions with constraints. For each of our axioms, we show that although it is challenging to satisfy these properties in general constrained instances, when one hones in on a restricted—yet highly expressive—class of constraints, we can achieve proportionality guarantees that represent approximations of desirable justified-representation axioms. In doing so, we also define novel adaptations of recently studied decision rules to our public-decision setting with constraints, namely the method of equal shares (MES) and the MeCorA rule. Finally, we adapt the priceability notion from the literature on committee elections, which provides another promising route to introduce proportionality into public decisions under constraints. A detailed summary of our results can be found in Table 1.
@@
This paper is structured as follows. We begin by discussing related work in Section 2. We then continue with Section 3 where we detail the constrained public-decision model, discuss the existing axioms of justified representation for public decisions, and also present the notion of deviating groups. Then each of Sections 4 and 5 deals with a particular public-decision interpretation of justified representation, namely via cohesiveness and via agreement. Before concluding in Section 7, we deal with our constrained version of the priceability axiom in Section 6.
2 Related Work
We begin by noting that our constrained public-decision model closely resembles that of judgment aggregation and it also naturally fits into the area of collective decisions in combinatorial domains (see Endriss, 2016, and Lang and Xia, 2016, for general introductions to these two topics, respectively).
Most relevant to our paper is the recent work conducted on fairness in the context of public decisions without constraints (Conitzer et al., 2017; Freeman et al., 2020; Skowron and Górecki, 2022). Conitzer et al. (2017) focused on individually proportional outcomes, thus, our work more closely aligns with that of Freeman et al. (2020) and Skowron and Górecki (2022) who import to public decisions the notion of justified representation (Aziz et al., 2017; Fernández et al., 2017; Peters and Skowron, 2020) from the literature on committee elections (Lackner and Skowron, 2023). Proportionality has also been studied in models of sequential decision-making that are relevant to our own as they can be seen as generalisations of the public-decision model without constraints (Bulteau et al., 2021; Chandak et al., 2024; Lackner, 2020). Amongst these sequential decision-making papers, those of Bulteau et al. (2021) and Chandak et al. (2024) relate to our work the most as they also implement justified-representation notions. In particular, Chandak et al. study a model similar to ours, where decisions involve multiple alternatives. The key distinction is that our work incorporates feasibility constraints on the final outcomes.
Proportionality under constrains have been studied extensively by Masarík et al. (2024) in a general social-choice model that can model both the unconstrained and constrained versions of public-decisions, along with a variety of other social choice domains. Their model elicits approval ballots from voters over a set of alternatives, and imposes a feasibility constraint on potential outcomes. The most notable differences with the model we work with are that $(1)$ while our proportionality axioms—based on the idea of feasible voter deviations—are structurally similar to those of Masarík et al., our axioms demand less from voter groups when evaluating their deserved representation, making them arguably more natural but also harder to satisfy; and $(2)$ by introducing properties tailored to our specific setting, we define and analyse constrained public-decision rules not considered by Masarík et al., such as the method of equal shares (an open question explicitly mentioned by Masarík et al.).
In a similar line, Mavrov et al. (2023) adapted justified representation for the committee elections model with arbitrary constraints, mostly focused on the notion of the core but also defining a version of extended justified representation. We refer to the discussion by Masarík et al. (2024) for illustrating the limitations of their approach.
In related fields, previous work studied proportionality in various models that feature collective choices on interconnected propositions: the belief merging setting (Haret et al., 2020), interdependent binary issues via conditional ballots (Brill et al., 2023), committee elections with synergies (Izsak et al., 2018), participatory budgeting with project interactions (Jain et al., 2020), and approval-based shortlisting with constraints, in a model of judgment aggregation by Chingoma et al. (2022).
Proportionality Satisfiability results
$\alpha\text{-}{\text{cohEJR}_{\mathcal{C}}}\text{-}\beta$ $\text{MES}_{\mathcal{C}}$ with $\lambda_{\text{unit}}$ for unconstrained elections with $\alpha=\nicefrac{{2}}{{(\max_{t\in[m]}|D_{t}|)}}$ and $\beta=1$ (Prop. 12)
$\Uparrow$
$\text{cohEJR}_{\mathcal{C}}$-1
$\text{MES}_{\mathcal{C}}$ with $\lambda_{\text{unit}}$ for binary unconstrained elections (Prop. 12)
for binary elections with $m=2$ (Prop. 10)
$\Uparrow$
$\text{cohEJR}_{\mathcal{C}}$
Not always satisfiable (Chandak et al., 2024)
when $|\mathcal{C}|=2$ and $\mathcal{C}$ has NFD (Prop. 4)
for binary elections with $m\leqslant 3$ and $\mathcal{C}$ has NFD (Prop. 5)
$\Downarrow$
$\text{cohJR}_{\mathcal{C}}$ Not always satisfiable (Prop. 6)
$\alpha\text{-}{\text{agrEJR}_{\mathcal{C}}}\text{-}\beta$
$\text{MES}_{\mathcal{C}}$ with $\lambda_{\text{fix}}$ for binary elections, $k$-restrictive $\mathcal{C}$ with $\alpha=\nicefrac{{1}}{{k}}$ and $\beta=1$ (Thm. 14)
Greedy $\text{MeCorA}_{\mathcal{C}}\text{-}(k-1)$ when $\mathcal{C}$-consistent, $k$-restrictive $\mathcal{C}$ (Thm. 15)
LS-$\text{PAV}_{\mathcal{C}}$ when $\mathcal{C}$-consistent, $k$-restrictive $\mathcal{C}$ with $\alpha=\nicefrac{{2}}{{(k+1)}}$ (Thm. 17)
$\Uparrow$
$\text{agrEJR}_{\mathcal{C}}$ Not always satisfiable even when $\mathcal{C}$ has NFD (Cor. 13)
Priceability $\text{MeCorA}_{\mathcal{C}}$ (Prop. 19)
Table 1: Table summarising our main results. The first column presents the constraints-adapted notions of proportionality we consider, and their relative implications. The second column presents results on their satisfiability: symbol identifies results where the axiom is satisfied by one of our proposed rules, and mentions the relevant theorem or reference.
@@ 3 The Model
A finite set of $n$ voters $N=\{1,\ldots,n\}$ has to take a collective decision on a finite set of $m$ issues $\mathcal{I}=\{a_{1},\ldots,a_{m}\}$. Typically, the public decision setting considers only two available decisions per issue, but we instead adopt the following more general setup. Taking $X$ as a set of alternatives, each issue $a_{t}\in\mathcal{I}$ is associated with its own finite set of alternatives called a domain $D_{t}=\{d_{t}^{1},d_{t}^{2},\ldots\}\subseteq X$ where $|D_{t}|\geqslant 2$ holds for all $t\in[m]$ with $X$. The design decision to go beyond binary issues is motivated by the wider real-life applicability of this model. Each voter $i\in N$ submits a ballot $\boldsymbol{b}_{i}=(\boldsymbol{b}_{i,1},\ldots,\boldsymbol{b}_{i,m})\in D_{1}\times\ldots\times D_{m}$ where an entry $\boldsymbol{b}_{i,t}$ indicates that voter $i$ chooses alternative $c$ for the issue $a_{t}$. A profile $\boldsymbol{B}=(\boldsymbol{b}_{1},\ldots,\boldsymbol{b}_{n})\in(D_{1}\times\ldots\times D_{m})^{n}$ is a vector of the $n$ voters’ ballots. An outcome $\boldsymbol{w}=(w_{1},\ldots,w_{m})\in D_{1}\times\ldots\times D_{m}$ is then a vector providing a decision for every issue at stake.
@@
We focus on situations where some constraints limit the set of possible collective outcomes: we denote by $\mathcal{C}\subseteq D_{1}\times\ldots\times D_{m}$ the non-empty set of feasible outcomes. We write $(\boldsymbol{B},\mathcal{C})$ to denote an election instance. By a slight abuse of notation we also refer to $\mathcal{C}$ as the constraint, and thus, we refer to elections instances where $\mathcal{C}=D_{1}\times\ldots\times D_{m}$ as unconstrained election instances. [^2] Note that while we work formally with the constraint being an enumeration of all feasible outcomes, in practice, it is often possible to represent the set of feasible outcomes in more concise forms—via the use of formulas of propositional logic, for example—to help with parsing said constraint and/or speed up computation by exploiting the constraint’s representation structure. Note that voter ballots need not be consistent with the constraints, i.e., for an election instance $(\boldsymbol{B},\mathcal{C})$, we do not require that $\boldsymbol{b}_{i}\in\mathcal{C}$ for all voters $i\in N$. [^3] This assumption takes our model closer to the particular model of judgment aggregation where the constraints on the output may differ from the constraints imposed on the voters’ input judgments (Endriss, 2018; Chingoma et al., 2022). This assumption is common in a number of settings of social choice. In committee elections, voters can approve more candidates than the committee target size while in participatory budgeting, the sum of the costs of a voter’s approved projects may exceed the instance’s budget. When deviating from this assumption we may explicitly require that voter ballots be constraint-consistent. For our setting, we argue that applying constraints only on the outcome of the collective decision helps capture real-world, constrained decision-making scenarios where either the constraint is uncertain when voters submit their ballots, or possibly, the voting process becomes more burdensome for voters as they attempt to create ballots with respect to a (possibly difficult to understand) constraint.
@@
A group of friends is deciding on the travel destinations of their shared holiday across France, visiting different regions. On a booking platform, there are a certain number of cities that can be selected per region such as: Toulouse, Carcassonne and Montpellier in Occitania; Rouen and Caen in Normandy; Rennes and Brest in Brittany; and so on. Each friend has a preferred combination of cities and their collective itinerary is subject to factors such as their travel budget or the available connections between cities. However, as costs and connections may change significantly on a day-to-day basis, it may be unclear which combination of cities are affordable. Therefore, it is not reasonable to impose the default requirement that voter ballots are constraint-consistent.
At times, we shall restrict to election instances where $D_{t}=\{0,1\}$ for every issue $a_{t}$. We refer to such cases as binary election instances. We explicitly state whether any result hinges on the restriction to binary instances. Given an outcome $\boldsymbol{w}$ for a binary instance, we define the vector $\boldsymbol{\bar{w}}=({\bar{w}}_{1},\ldots,{\bar{w}}_{m})$ as $\bar{w}_{t}=1-w_{t}$ for all issues $a_{t}\in\mathcal{I}$.
Let us now introduce some useful notation. Consider an outcome $\boldsymbol{w}$, a set of issues $S\subseteq\mathcal{I}$ and some vector $\boldsymbol{v}=(v_{1},\ldots,v_{m})\in D_{1}\times\ldots\times D_{m}$ (that can be interpreted as either an outcome or a voter’s ballot). We write ${\boldsymbol{w}}[{S\leftarrow\boldsymbol{v}}]=(w_{1}^{\prime},\ldots,w_{m}^{\prime})$ where $w_{t}^{\prime}=w_{t}$ for all issues $a_{t}\in\mathcal{I}\setminus{S}$ and $w_{t}^{\prime}=v_{t}$ for all issues $a_{j}\in S$. In other words, ${\boldsymbol{w}}[{S\leftarrow\boldsymbol{v}}]$ is the resultant vector of updating outcome $\boldsymbol{w}$’s decisions on the issues in $S$ by fixing them to those of vector $\boldsymbol{v}$. For a given issue $a_{t}\in\mathcal{I}$ and a decision $d\in D_{t}$, we use $N(a_{t},d)=\{i\in N\mid\boldsymbol{b}_{i,t}=d\}$ to denote the set of voters that agree with decision $d$ on issue $a_{t}$. Given two vectors $\boldsymbol{v},\boldsymbol{v}^{\prime}\in D_{1}\times\ldots\times D_{m}$, we denote the agreement between them by $\text{\rm Agr}(\boldsymbol{v},\boldsymbol{v}^{\prime})=\{a_{t}\in\mathcal{I}\mid v_{t}={v}_{t}^{\prime}\}$.
We define the satisfaction that a voter $i$ obtains from an outcome $\boldsymbol{w}$ as $u_{i}(\boldsymbol{w})=|\text{\rm Agr}(\boldsymbol{b}_{i},\boldsymbol{w})|$, i.e., the number of decisions on which the voter $i$ is in agreement with outcome $\boldsymbol{w}$. We adopted this notion of satisfaction as it is a common choice within the literature on public decisions (Freeman et al., 2020; Skowron and Górecki, 2022), which is in turn grounded on a similar assumption from work on committee elections (Lackner and Skowron, 2023).
3.1 Proportionality via Justified Representation
This section starts with the observation that classical notions of proportionality fall short when considering interconnected decisions (see Example 2), and then follows with our adaptations of such axioms for constrained environments.
Ideally, when looking to make a proportional collective choice, we would like to meet the following criteria: a group of similarly-minded voters that is an $\alpha$-fraction of the population should have their opinions reflected in an $\alpha$-fraction of the $m$ issues. In the setting of committee elections, this is formally captured via the justified representation axioms, with one of the most widely studied being extended justified representation (EJR) (Aziz et al., 2017). In the setting of public decisions there are two adaptations that have been studied. The first approach intuitively states that “a group of voters that agree on an $\alpha$-fraction of the issues in $\mathcal{I}$ and are $\alpha$-fraction of the voter population, should have some control over an $\alpha\cdot|\mathcal{I}|$ number of the issues in $\mathcal{I}$(Chandak et al., 2024; Freeman et al., 2020). In this approach, the requirements on the voter groups are captured by the notion of cohesiveness and so we refer to this version of EJR as cohesiveness-EJR. The second approach intuitively states that “a group of voters that agree on a set of issues $T$ and represent an $\alpha$-fraction of the voter population, should have some control over an $\alpha\cdot|T|$ number of the total issues in $\mathcal{I}$(Chandak et al., 2024; Masarík et al., 2024; Skowron and Górecki, 2022). We refer to it as agreement-EJR. Observe that cohesiveness-EJR is stronger than, and implies, agreement-EJR.
Meeting the ideal outlined by both of these notions is not easy in our setting, as the constraint $\mathcal{C}$ could rule out a seemingly fair outcome from the onset.
Example 2.